One integer is 7 more than another. Their product is 294. Find the integers.
Answer by stanbon(75887)

One integer is 7 more than another. Their product is 294. Find the integers.
------
smaller:: x
larger:: x + 7
------
Equation:
x(x+7) = 294
x^2 + 7x - 294 = 0
(x-14)(x+21) = 0
------
x = 14
x+7 = 21
--------------
Cheers,
Stan H.
-----------